OpenStudy (anonymous):

Solve the equation. |10x – 14| - 5 = 156

OpenStudy (anonymous):

|10x – 14| - 5 = 156 |10x – 14| = 156 +5

OpenStudy (anonymous):

and so on...

OpenStudy (anonymous):

so 10x - 14 = 161 10x - 14 = -161

OpenStudy (anonymous):

SO WHAT WOULD THE ANSWER BE?